A plano convex lens placed in air has curved surface aving radius of radius of curvature 60cm and it is made of refractive index 1.5.When the convex surface is silvered, the system will work as a concave mirror of focal length?

Grade:12

1 Answers

Shyama P
12 years ago

(1/f)=(1/fmirror)+(2/flens)

1/fmirror=2/R

1/flens=(1.5-1)/R   lens makers formula

         =1/2R

Sub.

1/f=3/R => f=20cm